# A Theory of Recommendations

### v1

@@ A Theory of Recommendations
Jean-Michel Benkert and Armin Schmutzler Benkert: Department of Economics, University of Bern, jean-michel.benkert@unibe.ch. Schmutzler: Department of Economics, University of Zurich, and CEPR, armin.schmutzler@econ.uzh.ch. We are grateful for very helpful discussions with Piotr Dworczak, Igor Letina, Marc Möller, Nick Netzer, Volker Nocke, Egor Starkov, Jakub Steiner and seminar participants in Bern, Fribourg, Zurich, at SAET (2023), Swiss Society of Economics of Statistics (2024), and Swiss IO Day (2024).
(August 21, 2024)
@@ 1 Introduction
These features will be even more salient for ratings provided on internet platforms, as such ratings usually do not come from close acquaintances. Hence, the receivers should consider that senders have preferences which differ from their own and are likely to give an incomplete account of their consumption experience without exerting much effort. Backing out the implications of the recommendation for the receiver’s own decision is therefore a complex task. Relevant considerations for the interpretation are these: To what extent does the recommendation contain objective truth? And when it does not, how likely is it that the preferences of the sender and receiver are aligned? The answers to these questions will determine whether it is in the receiver’s interest to follow the recommendation and how valuable it is.
Our model captures the above three features of recommendations in the simplest possible way. An agent must decide between several options, referred to as products for ease of exposition. Products are two-dimensional objects, differing only in whether they have high or low quality in each dimension. Thus, four product versions exist. All agents agree that a product is worth choosing if it has high quality in both dimensions and is thus (objectively) good. Conversely, they agree that a product is not worth choosing if it has low quality in both dimensions and is thus (objectively) bad. Accordingly, we call the products uncontroversial in both of these cases. By contrast, we refer to products as controversial when they have high quality in one dimension and low quality in the other. An agent prefers a controversial product that has high quality in the dimension that is more relevant for them, and is thus subjectively good, to a product that has high quality only in the less-relevant dimension (thus subjectively bad). Moreover, even agents who agree about which controversial product they prefer may differ in the intensity of these preferences. This heterogeneity is captured in a continuous type distribution.
@@
To answer these questions, we need to derive intermediate results of independent interest. Most importantly, we investigate under which conditions receivers accept a recommendation. We find that for a Bayesian expected-payoff maximizer, two key quantities measuring the objective and subjective content of the recommendation are decisive for receiver behavior. These quantities are simple functions of the model primitives. Intuitively, the objective informational content of a recommendation is high if it makes the product sufficiently more likely to be good and less likely to be bad. The subjective content captures the effect of the recommendation on the relative probability of the two controversial product versions. The evaluation of the subjective content thus depends on the receiver’s type. When the recommendation bears little or no subjective content, all types accept it, thus buying after a buy recommendation and choosing the outside option (one of the other alternatives) after a don’t-buy recommendation. As the subjective content increases, revealing that one of the two controversial product versions becomes more likely, only receivers without strong preferences for the other version accept the recommendation. When the recommendation threshold $R$ approaches the maximal payoff, a buy recommendation essentially becomes objective, and similarly for a don’t-buy recommendation when $R$ approaches the minimal payoff. As receivers are aware of this, they accept the recommendation in each of these polar cases since the subjective component vanishes.
@@
Based on this analysis, we then study the value of recommendation systems. We think of this value as the expected-payoff increase of a randomly chosen receiver resulting from a recommendation given by a randomly chosen sender. When the recommendation is objective, or the subjective content is small, the analysis is comparatively simple: As all receivers accept the recommendation, they choose the outside option after a don’t-buy recommendation, just as if they had not received any recommendation at all. Therefore, only buy recommendations create value—they do so by inducing the receiver to purchase the product if it generates a higher expected payoff for him than the outside option. When the recommendation contains more subjective content, only a subset of the receivers accept it, rendering the analysis more complex. The value of the recommendation system then has two sources. First, receivers who accept buy recommendations benefit because they expect the recommended product to be better than the outside option. Second, receivers who do not accept the recommendation benefit from buying the product in spite of a don’t-buy recommendation: These receivers have preferences that are not well aligned with the general population so that the don’t-buy recommendation provides good news about the chances of obtaining a subjectively good product rather than one they like less.
@@
Next, we turn to the design of recommendation systems. Motivated by the practice of digital platforms, we consider the recommendation threshold $R$ as a design variable. When the type distribution is symmetric, so that there is no preference bias on average, the value of the recommendation system is a monotone function of the threshold. If the value is increasing, the product optimally receives a buy recommendation only if it is good; if the value is decreasing, the product optimally receives a don’t-buy recommendation only if it is bad. The former case arises if good products are more likely ex ante than bad products are, and conversely for the latter case. Either way, the content of the recommendation essentially becomes objective. In contrast, with an asymmetrically distributed population, value maximization does not necessarily require an extreme threshold. Rather than revealing only whether a product is good or bad, it may be optimal for the sender to give a buy recommendation even if the product is controversial but sufficiently aligned with her preferences. Intuitively, giving up the exclusive focus on objective content makes sense when both controversial product versions are common but there is broad agreement on which of them is preferable. By contrast, extreme recommendation thresholds that only reveal objective content are still desirable if the product is much more likely to be objectively good than objectively bad or vice versa.
@@
Finally, we show the robustness of our results. With mild adjustments, the main insights continue to hold when sender and receiver distributions are distinct—a situation that could arise because, for instance, recommendations stem from early adopters who do not represent the population at large. Second, we allow for more than two types of recommendations. Again, extreme recommendation thresholds increase the probability that a recommendation has objective content and thus increase the value of the system when the population is symmetrically distributed. Third, we show that having access to many recommendations does not necessarily change receiver behavior compared to the case of a single recommendation with an optimal threshold. Then, having access to multiple recommendations does not increase the value.
Section 2 relates the paper to the literature. Section 3 introduces the model and discusses its key assumptions. In Section 4, we describe the conditions under which different receiver types accept recommendations. Then, in Section 5 we first characterize the value of a recommendation system before investigating the determinants of the value and the design of value-maximizing systems. Section 6 discusses our results and considers extensions. Section 7 concludes. All proofs are relegated to the appendix.
@@ 3.1 Assumptions
We consider a consumer’s choice of a product from a large set of alternatives. We assume that each product is fully characterized by a two-dimensional quality vector, where $Q_{d}=1$ and $Q_{d}=0$ capture high and low quality, respectively, in dimension $d\in\{1,2\}$. Accordingly, four different product versions exist. We assume that the products are experience goods. Thus, the consumer cannot observe product quality before consumption, so the alternatives are identical for him ex ante. However, he knows the prior distribution of the quality vectors, referred to as quality distribution for brevity. Consumers have heterogeneous preferences fully characterized by their type $i\in[-1/2,1/2]$. A consumer of type $i$ receives a payoff of
@@
$\pi^{B}(R)$ $:=q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R),$ (1)
$\pi^{D}(R)$ $:=q_{1}(1-\phi_{1}(R))+q_{2}(1-\phi_{2}(R))+q_{L},$ (2)
where $\pi^{D}(R)=1-\pi^{B}(R)$, and $\phi_{1}(R):=1-F(R-1/2)$ and $\phi_{2}(R):=F(1/2-R)$ denote the probability of a buy recommendation conditional on the product version being $(1,0)$ or $(0,1)$, respectively. In words, the probability of receiving a buy recommendation from a randomly chosen sender is given by the sum of the probability that a product is objectively good and the probability that it is subjectively good enough from a randomly chosen sender’s perspective. Thus, following a buy recommendation the receiver’s posterior $\mathbf{p}^{B}(R):=(p_{H}^{B}(R),p_{1}^{B}(R),p_{2}^{B}(R),p_{L}^{B}(R))$ becomes
@@ 3.2 Discussion of the Framework
With a single threshold above and below which senders give buy and don’t-buy recommendations, respectively, it is impossible to always reveal the product version. Thus, as desired, recommendations provide coarse information. If, in contrast, the sender could give four different ratings depending on the consumption experience, full revelation would be possible in our parsimonious model with only four product versions. However, in practice the greater complexity of most products would typically render such full revelation infeasible.
@@ 4 Optimal Receiver Behavior
In this section, we ask under which circumstances a receiver will accept recommendations, depending on his type $i$ and the recommendation system $\mathcal{R}$. In Section 4.1, we discuss through which channels recommendations affect beliefs, distinguishing between the objective and subjective content of the recommendations. Section 4.2 then characterizes the optimal behavior of the receiver by identifying thresholds which divide the type set into receivers that accept the recommendation and those that do not.
@@ 4.1 The Effects of Recommendations on Beliefs
Recommendations contain coarse information about the product under consideration, leading the receiver to update his beliefs about its quality. We first focus on buy recommendations, which have the following intuitive effects: [^11] In Appendix A.1, we provide a more formal analysis, decomposing the difference between posteriors and priors into three corresponding components.
@@
The first two effects in this belief-updating decomposition are objectively positive: They correspond to increases of the expected payoff for every receiver type, albeit of different sizes. By contrast, the third effect is fully type-dependent and thus subjective, resulting in an increase in the payoff for some receivers and a decrease for others. The following distinction is useful:
Definition 2.
A recommendation $r\in\{B,D\}$ that induces posteriors $\mathbf{p}^{r}$ given prior $\mathbf{q}$
  1. (i)
    shifts probability from $(0,1)$ to $(1,0)$ if $p_{2}^{r}-q_{2}<p_{1}^{r}-q_{1}$;
  2. (ii)
    shifts probability from $(1,0)$ to $(0,1)$ if $p_{2}^{r}-q_{2}>p_{1}^{r}-q_{1}$;
  3. (iii)
    does not affect the probability between $(0,1)$ and $(1,0)$ otherwise.
As we will lay out below, whether a receiver considers the subjective content of a recommendation as positive or negative depends on his type $i$ and on the direction in which the recommendation shifts probability.
@@ 4.2 Characterizing Optimal Receiver Behavior
We are now in a position to understand whether it is optimal for the receiver to accept a recommendation. To this end, denote by $U_{i}^{B}$ and $U_{i}^{D}$ player $i$’s expected payoff of buying a product with a buy and don’t-buy recommendation, respectively, and by $U_{i}^{0}$ the expected payoff of buying the alternative product with no recommendation. We have
$U_{i}^{r}$ $:=p_{H}^{r}+(1/2+i)p_{1}^{r}+(1/2-i)p_{2}^{r},\quad r\in\{B,D\};$ (3)
$U_{i}^{0}$ $:=q_{H}+(1/2+i)q_{1}+(1/2-i)q_{2}.$ (4)
For an expected-payoff maximizer, it is optimal to accept a buy recommendation if and only if $U_{i}^{B}-U_{i}^{0}\geq 0$, that is, the expected payoff from buying the recommended product is at least as high as the expected payoff from buying one of the alternative products (that have no recommendation). [^12] We use the tie-breaking rule that indifferent agents accept the recommendation. This condition can equivalently be written as
@@
To formulate this in compact form, we introduce the following terminology.
@@
We can now restate condition (5) as
@@
We use this formulation to justify our terminology. For the unbiased receiver $i=0$, equation (8) simplifies to $\Delta_{O}^{B}\geq 0$. As this receiver does not care about the shift in probabilities between the two controversial versions of the products, $\Delta_{O}^{B}\geq 0$ fully reflects the two objective effects (I) and (II) in the belief-updating decomposition in Section 4.1. For $i\neq 0$, the issue is more subtle. As long as $q_{1}=q_{2}$, the expected payoff conditional on the information that a product is controversial is $1/2$ for all $i$ and thus type-independent. Therefore, all types gain the same amount from (I) learning that a product is good or controversial instead of bad; the same applies to (II) learning that a product is good instead of controversial. For $q_{1}=q_{2}$, $\Delta_{O}^{B}$ therefore exactly captures these two type-independent (and thus objective) effects of a buy recommendation. [^13] The value of learning that the recommended product is more likely to be good is $p_{H}^{r}-q_{H}$, and the value of the change in the probability that it is controversial is $\frac{p_{1}^{r}+p_{2}^{r}-(q_{1}+q_{2})}{2}$. For $i\neq 0$ and $q_{1}\neq q_{2}$, the issue is complicated further. Though the effects captured by (I) and (II) in Section 4.1 are positive for all types, their size depends on the receiver’s type. Intuitively, if one of the controversial product versions (say (1,0))(1,0))( 1 , 0 ) ) is more likely than the other ex ante, learning that a product is controversial rather than bad is more valuable for types who prefer $(1,0)$ over $(0,1)$. Similarly, learning that the product is good rather than controversial is less valuable for those types. Even in this case, the interpretation of $\Delta_{O}^{B}\geq 0$ as an objective effect is justified because of the general agreement on its sign. We can now state our first result.
@@
    accepts a buy recommendation if and only if he accepts a don’t-buy recommendation;
@@
    accepts a recommendation if and only if $\Delta_{O}^{B}\geq i\Delta_{S}^{B}$.
Intuitively, a receiver accepts a recommendation if he thinks that it has sufficient objective content and/or that his preferences are sufficiently aligned with the sender’s. Whether the recommendation is positive or negative does not matter for this assessment. Rather, as laid out in the decomposition in Section 4.1, the objective effect $\Delta_{O}^{B}$ of a buy recommendation is positive, as the latter makes a good product more likely relative to the controversial products and rules out a bad one. By contrast, the sign of the subjective effect $\Delta_{S}^{B}$ can be positive or negative. In short, a receiver accepts a buy recommendation unless the subjective effect is sufficiently large and his preferences are not sufficiently well aligned with those of the rest of the population.
Denoting the type who is indifferent between accepting and not accepting recommendations as
$\tilde{i}:=$ $\frac{\Delta_{O}^{B}}{\Delta_{S}^{B}}\text{ if }\Delta_{S}^{B}\neq 0\text{,}$ (9)
@@
    If $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, then all $i\in[-1/2,1/2]$ accept the recommendation.
@@
    If $\Delta_{S}^{B}<-2\Delta_{O}^{B}$, then all $i\geq\tilde{i}$ accept the recommendation.
@@
    If $\Delta_{S}^{B}>2\Delta_{O}^{B}$, then all $i\leq\tilde{i}$ accept the recommendation.
The intuition for Corollary 1, the proof of which is omitted, is straightforward. For $\Delta_{S}^{B}=0$, the recommendation does not shift probability between the controversial product versions. Accordingly, accepting the recommendation is optimal for all players because of its positive objective effect $\Delta_{O}^{B}\geq 0$. For $\Delta_{S}^{B}\neq 0$, the conclusion is the same as long as $|\Delta_{S}^{B}|$ remains small enough—as required by (i). In that case, even the receivers who are most adversely affected by the subjective content of the recommendation (those at $i=1/2$ or $i=-1/2$) are convinced by its relatively objective character. As $|\Delta_{S}^{B}|$ increases, the role of the probability shift becomes apparent. As types $i>0$ prefer (1,0) over (0,1), a shift of the probability from the latter to the former induced by a recommendation ($\Delta_{S}^{B}<0$) increases such receivers’ expected payoff, reinforcing the objective effect. By contrast, such a shift reduces the expected payoff of types $i<0$. The condition in (ii) always holds for $i>0$, as the objective and subjective effects reinforce each other. It also holds for $i<0$ as long as $|i|<|\tilde{i}|$ so that the adverse subjective effect does not dominate the objective effect. Similarly, when the recommendation shifts the probability from (1,0) to (0,1) (when $\Delta_{S}^{B}>0$), all types $i<0$ accept. By contrast, types $i>0$ experience a reduction in their value, which can outweigh the positive objective effect. The condition in (iii) makes sure that the latter case does not arise. [^14] The effect on the expected payoffs in cases (ii) and (iii) is similar to the effect of a demand rotation as discussed in Johnson and Myatt (2006): For the types who accept the recommendation, the expected payoffs increase; for those who do not, they fall.
@@
Suppose $R\rightarrow 1$ or $R\rightarrow 0$. Then, all types accept the recommendation irrespective of the product distribution.
In the limits, both types of recommendations essentially contain only objective information. For instance, when $R\rightarrow 1$, buy recommendations require that the product is objectively good, while don’t-buy recommendations require that it is not. Either way, the recommendation does not reveal anything about the relative probabilities of the two controversial product versions. Instead, it provides objective information that is equally valuable for all types, irrespective of their preferences for controversial products. Thus, all types accept the recommendation.

### v2

@@ A Theory of Recommendations
Jean-Michel Benkert and Armin Schmutzler Benkert: Department of Economics, University of Bern, jean-michel.benkert@unibe.ch. Schmutzler: Department of Economics, University of Zurich, and CEPR, armin.schmutzler@econ.uzh.ch. We are grateful for very helpful discussions with Piotr Dworczak, Johannes Johnen, Igor Letina, Marc Möller, Nick Netzer, Volker Nocke, Egor Starkov, Jakub Steiner and seminar participants in Bern, Fribourg, Zurich, at SAET (2023), Swiss Society of Economics of Statistics (2024), Swiss IO Day (2024) and Verein für Socialpolitik (2024).
(October 11, 2024)
@@ 1 Introduction
These features will be even more salient for ratings provided on internet platforms, as such ratings usually do not come from close acquaintances. Hence, the receivers should consider that senders have preferences which differ from their own and are likely to give an incomplete account of their consumption experience without exerting much effort. Backing out the implications of the recommendation for the receiver’s own decision is therefore a complex task. Relevant considerations for the interpretation are these: To what extent does the recommendation contain objective truth? And when it does not, how likely is it that the preferences of the sender and receiver are aligned? The answers to these questions will determine whether it is in the receiver’s interest to accept the recommendation and how valuable it is.
Our model captures the above three features of recommendations in the simplest possible way. An agent must decide between several options, referred to as products for ease of exposition. Products are two-dimensional objects, differing only in whether they have high or low quality in each dimension. Thus, four product versions exist. There is a continuum of agent types. All types agree that a product is worth choosing if it has high quality in both dimensions and is thus (objectively) good. Conversely, they agree that a product is not worth choosing if it has low quality in both dimensions and is thus (objectively) bad. Accordingly, we call the products uncontroversial in both of these cases. By contrast, we refer to products as controversial when they have high quality in one dimension and low quality in the other. An agent prefers a controversial product that has high quality in the dimension that is more relevant for them, and is thus subjectively good, to a product that has high quality only in the less-relevant dimension (thus subjectively bad). Moreover, even agents who agree about which controversial product they prefer may differ in the intensity of these preferences (as captured by the continuous type distribution).
@@
To answer these questions, we need to derive intermediate results of independent interest. Most importantly, we investigate under which conditions receivers accept all recommendations. We find that for a Bayesian expected-payoff maximizer, two key quantities measuring the objective and subjective content of the recommendation are decisive for receiver behavior. These quantities are simple functions of the model primitives. Intuitively, the objective informational content of a recommendation is high if it makes the product sufficiently more likely to be (objectively) good and less likely to be bad. The subjective content captures the effect of the recommendation on the relative probability of the two controversial product versions. The evaluation of the subjective content thus depends on the receiver’s type. When the recommendation bears little or no subjective content, all types accept all recommendations, thus buying after a buy recommendation and choosing the outside option (one of the other alternatives) after a don’t-buy recommendation. As the subjective content increases, revealing that one of the two controversial product versions becomes more likely, only receivers without strong preferences for the other version accept all recommendations. When the recommendation threshold $R$ approaches the maximal payoff, a buy recommendation essentially becomes objective, and similarly for a don’t-buy recommendation when $R$ approaches the minimal payoff. As receivers are aware of this, they accept all recommendations in each of these polar cases since the subjective component vanishes.
@@
Based on this analysis, we then study the value of recommendation systems. We think of this value as the expected-payoff increase of a randomly chosen receiver resulting from a recommendation given by a randomly chosen sender. When the recommendation is objective, or the subjective content is small, the analysis is comparatively simple: As all receivers accept all recommendations, they choose the outside option after a don’t-buy recommendation, just as if they had not received any recommendation at all. Therefore, only buy recommendations create value—they do so by inducing the receiver to purchase the product if it generates a higher expected payoff for him than the outside option. When the recommendation contains more subjective content, only a subset of the receivers accept it, rendering the analysis more complex. The value of the recommendation system then has two sources. First, receivers who accept all recommendations benefit because they expect the recommended product to be better than the outside option. Second, receivers who never accept recommendations benefit from buying the product in spite of a don’t-buy recommendation: These receivers have preferences that are not well aligned with the general population so that the don’t-buy recommendation provides good news about the chances of obtaining a subjectively good product rather than one they like less.
@@
Next, we turn to the design of recommendation systems. Motivated by the practice of digital platforms, we consider the recommendation threshold $R$ as a design variable. When the type distribution is symmetric, so that there is no preference bias on average, the value of the recommendation system is a monotone function of the threshold. If the value is increasing, the product optimally receives a buy recommendation only if it is good; if the value is decreasing, the product optimally receives a don’t-buy recommendation only if it is bad. The former case arises if good products are more likely ex ante than bad products are, and conversely for the latter case. Either way, the content of the recommendation essentially becomes objective. In contrast, with an asymmetrically distributed population, value maximization does not necessarily require an extreme threshold. Rather than revealing only whether a product is good or bad, it may be optimal for the sender to give a buy recommendation even if the product is controversial but sufficiently aligned with her preferences. Intuitively, giving up the exclusive focus on objective content makes sense when both controversial product versions are common and there is broad agreement on which of them is preferable. By contrast, extreme recommendation thresholds that only reveal objective content are still desirable if the product is much more likely to be objectively good than objectively bad or vice versa.
@@
Finally, we show the robustness of our results. With mild adjustments, the main insights continue to hold when sender and receiver distributions are distinct—a situation that could arise because, for instance, recommendations stem from early adopters who do not represent the population at large. This also has non-trivial design implications: Our results on the relation between mean-preserving spreads can be adapted to show how a designer would want to choose the sender distribution, which may be at least partly under his control. Second, we allow for more than two types of recommendations. Again, extreme recommendation thresholds increase the probability that a recommendation has objective content and thus increase the value of the system when the population is symmetrically distributed. Third, we show that having access to many recommendations does not necessarily change receiver behavior compared to the case of a single recommendation with an optimal threshold. Then, having access to multiple recommendations does not increase the value.
Section 2 relates the paper to the literature. Section 3 introduces the model and discusses its key assumptions. In Section 4, we describe the conditions under which different receiver types accept all recommendations. Then, in Section 5 we first characterize the value of a recommendation system before investigating the determinants of the value and the design of value-maximizing systems. Section 6 discusses our results and considers extensions. Section 7 concludes. All proofs are relegated to the appendix.
@@ 3.1 Assumptions
We consider a consumer’s choice of a product from a set of alternatives $N\in\mathbb{N}$ with $N>1$. We assume that each product is fully characterized by a two-dimensional quality vector, where $Q_{d}=1$ and $Q_{d}=0$ capture high and low quality, respectively, in dimension $d\in\{1,2\}$. Accordingly, four different product versions exist. We assume that the products are experience goods. Thus, the consumer cannot observe product quality before consumption, so the alternatives are identical for him ex ante. However, he knows the prior distribution of the quality vectors, referred to as quality distribution for brevity. Consumers have heterogeneous preferences fully characterized by their type $i\in[-1/2,1/2]$. A consumer of type $i$ receives a payoff of
@@
$\pi^{B}(R)$ $:=q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R),$
$\pi^{D}(R)$ $:=q_{1}(1-\phi_{1}(R))+q_{2}(1-\phi_{2}(R))+q_{L},$
where $\pi^{D}(R)=1-\pi^{B}(R)$, and $\phi_{1}(R):=1-F(R-1/2)$ and $\phi_{2}(R):=F(1/2-R)$ denote the probability of a buy recommendation conditional on the product version being $(1,0)$ or $(0,1)$, respectively. In words, the probability of receiving a buy recommendation from a randomly chosen sender is given by the sum of the probability that a product is objectively good and the probability that it is subjectively good enough from a randomly chosen sender’s perspective. Thus, accepting a buy recommendation the receiver’s posterior $\mathbf{p}^{B}(R):=(p_{H}^{B}(R),p_{1}^{B}(R),p_{2}^{B}(R),p_{L}^{B}(R))$ becomes
@@ 3.2 Discussion of the Framework
With a single threshold above and below which senders give buy and don’t-buy recommendations, respectively, it is impossible to always reveal the product version. Thus, as desired, recommendations provide coarse information. If, in contrast, the sender could give four different ratings depending on the consumption experience, full revelation would be possible in our parsimonious model with only four product versions. In reality, recommendations can convey more content than whether a buyer liked a product or not, for instance, by including verbal descriptions. However, given the typical complexity of products, even relatively detailed recommendations will usually only transmit coarse information.
@@ 4 Optimal Receiver Behavior
In this section, we ask under which circumstances a receiver will accept all recommendations, depending on his type $i$ and the recommendation system $\mathcal{R}$, as introduced in Definition 1. In Section 4.1, we discuss through which channels recommendations affect beliefs, distinguishing between the objective and subjective content of the recommendations. Section 4.2 then characterizes the optimal behavior of the receiver by identifying thresholds which divide the type set into receivers that accept all recommendations and those that never do.
@@ 4.1 The Effects of Recommendations on Beliefs
Recommendations contain coarse information about the product under consideration, leading the receiver to update his beliefs about its quality. We first focus on buy recommendations, the total effect of which can be decomposed as follows: [^11] In Appendix A.1, we provide a more formal analysis, decomposing the difference between posteriors and priors into three corresponding components.
@@
The first two effects in this belief-updating decomposition are objectively positive: They correspond to increases of the expected payoff for every receiver type, albeit of different sizes. By contrast, the third effect is fully type-dependent and thus subjective, resulting in an increase in the payoff for some receivers and a decrease for others.
@@ 4.2 Characterizing Optimal Receiver Behavior
We are now in a position to understand whether it is optimal for the receiver to accept recommendations. To this end, note that, after receiving a recommendation, the receiver can either buy the recommended product or not buy it. Denote by $U^{B}(i)$ and $U^{D}(i)$ player $i$’s expected payoff of buying a product with a buy and don’t-buy recommendation, respectively, and by $U^{0}(i)$ the expected payoff of not buying, i.e., of buying the alternative product with no recommendation. We have
$U^{r}(i)$ $:=p_{H}^{r}+(1/2+i)p_{1}^{r}+(1/2-i)p_{2}^{r},\quad r\in\{B,D\};$
$U^{0}(i)$ $:=q_{H}+(1/2+i)q_{1}+(1/2-i)q_{2}.$
Now, the difference between $U^{r}(i)$ and $U^{0}(i)$ determines the behavior of an expected-payoff maximizing receiver as follows.
Definition 2.
We say that a receiver $i$
  1. (i)
    accepts a buy recommendation if $U^{B}(i)\geq U^{0}(i)$; otherwise we say that he does not accept a buy recommendation;
  2. (ii)
    accepts a don’t-buy recommendation if $U^{0}(i)\geq U^{D}(i)$; otherwise we say that he does not accept a don’t-buy recommendation
  3. (iii)
    accepts all recommendations if he accepts both buy and don’t-buy recommendations;
  4. (iv)
    never accepts recommendations if he does not accept either buy and don’t-buy recommendations.
Hence, a receiver who accepts all recommendations “does as he is told”, whereas a receiver who never accepts recommendations is a “contrarian” who does the opposite of what he is told. [^12] The use of the term aligns with Drehmann et al. (2005), who document contrarian behavior, i.e., investors going against the behavior of the predecessors, in an financial-markets experiment. Note that the above definition does not cover all conceivable types of behavior, as a receiver $i$ could accept buy recommendations and not accept don’t-buy recommendations, which is tantamount to always buying the recommended product and thus ignoring the content of the recommendations. Analogously, a receiver $i$ could ignore the content of the recommendations and never buy the recommended product. We will return to these cases shortly.
The condition from Definition 2(i) for receiver $i$ to accept a buy recommendation can equivalently be written as
@@
For more compact formulations, we introduce the following terminology.
@@
We can now restate condition (1) about accepting a buy recommendation as
@@
[…]
@@
    either accepts all recommendations or never accepts recommendations;
@@
    accepts all recommendations if and only if $\Delta_{O}^{B}\geq i\Delta_{S}^{B}$.
The first part of the result states that Definition 2 indeed captures all relevant forms of behavior. That is, no type of receiver will ignore the informational content of recommendations by always or never buying the recommended product. Intuitively, a receiver accepts all recommendations if he thinks that they have sufficient objective content and/or that his preferences are sufficiently aligned with the sender’s. Conversely, a receiver never accepts recommendations if he thinks that they do not have sufficient objective content and/or that his preferences are not sufficiently aligned with the sender’s. Whether the recommendation is positive or negative does not matter for this assessment. Rather, as laid out in the decomposition in Section 4.1, the objective effect $\Delta_{O}^{B}$ of a buy recommendation is positive, as the latter makes a good product more likely relative to the controversial products and rules out a bad one. By contrast, the sign of the subjective effect $\Delta_{S}^{B}$ can be positive or negative. The second part of the result shows which types accept all recommendations, which happens unless the subjective effect is sufficiently large and the receiver’s preferences are not sufficiently well aligned with those of the rest of the population. We emphasize that receivers who never accept recommendations still make use of them, by doing the opposite of what they are told.
Proposition 1 tells us that we can decompose the population into at most two different groups: those who accept all and those who never accept recommendations. Denoting the type who is indifferent between always accepting and never accepting recommendations as
$\tilde{i}:=$ $\frac{\Delta_{O}^{B}}{\Delta_{S}^{B}}\text{ if }\Delta_{S}^{B}\neq 0\text{,}$
@@
    If $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, then all $i\in[-1/2,1/2]$ accept all recommendations.
@@
    If $\Delta_{S}^{B}<-2\Delta_{O}^{B}$, then all $i\geq\tilde{i}$ accept all recommendations.
@@
    If $\Delta_{S}^{B}>2\Delta_{O}^{B}$, then all $i\leq\tilde{i}$ accept all recommendations.
The intuition for Corollary 1, the proof of which is omitted, is straightforward. Without loss, consider a buy recommendation. For $\Delta_{S}^{B}=0$, the recommendation does not shift probability between the controversial product versions. Accordingly, accepting the recommendation is optimal for all players because of its positive objective effect $\Delta_{O}^{B}\geq 0$. For $\Delta_{S}^{B}\neq 0$, the conclusion is the same as long as $|\Delta_{S}^{B}|$ remains small enough—as required by (i). In that case, even the receivers who are most adversely affected by the subjective content of the recommendation (those at $i=1/2$ or $i=-1/2$) are convinced by its relatively objective character. As $|\Delta_{S}^{B}|$ increases, the role of the probability shift becomes apparent. As types $i>0$ prefer (1,0) over (0,1), a shift of the probability from the latter to the former induced by a recommendation ($\Delta_{S}^{B}<0$) increases such receivers’ expected payoff, reinforcing the objective effect. By contrast, such a shift reduces the expected payoff of types $i<0$. The condition in (ii) always holds for $i>0$, as the objective and subjective effects reinforce each other. It also holds for $i<0$ as long as $|i|<|\tilde{i}|$ so that the adverse subjective effect does not dominate the objective effect. Similarly, when the recommendation shifts the probability from (1,0) to (0,1) (when $\Delta_{S}^{B}>0$), all types $i<0$ accept. By contrast, types $i>0$ experience a reduction in their value, which can outweigh the positive objective effect. The condition in (iii) makes sure that the latter case does not arise. [^15] The effect on the expected payoffs in cases (ii) and (iii) is similar to the effect of a demand rotation as discussed in Johnson and Myatt (2006): For the types who accept the recommendation, the expected payoffs increase; for those who do not, they fall.
Corollary 1 may potentially be useful to draw inferences about the relative importance of the objective and subjective effects of recommendations: If an observer sees a large share of receivers accepting recommendations, this can be interpreted as evidence for a relatively high objective effect of the recommendation.
@@
Suppose $R\rightarrow 1$ or $R\rightarrow 0$. Then, all types accept all recommendations irrespective of the product distribution.
In the limits, both types of recommendations essentially contain only objective information. For instance, when $R\rightarrow 1$, buy recommendations require that the product is objectively good, while don’t-buy recommendations require that it is not. Either way, the recommendation does not reveal anything about the relative probabilities of the two controversial product versions. Instead, it provides objective information that is equally valuable for all types, irrespective of their preferences for controversial products. Thus, all types accept all recommendations.